A ballon is filled to a volume of 1.5L with 3 moles of gas at 25 degrees celecius. With pressure and temperature help constant, what will be the volume (in L) of the ballon if .2 moles of gas are added


1
Expert's answer
2021-05-12T06:28:41-0400

Given moles of a gas =3mol

Volume of baloon=1.5L

Now when we add 0.2moles of the gas then the total moles of the gas becomes 3.2 mol

Suppose the new Volume of the gas =vL

"\therefore" Given that temperature and pressure are at constant then :

"=V\alpha n"

But:"\frac{V}{n}=\frac{V}{n}"


"\frac{1.5}{3}=\frac{V}{3.2}"


"V=\frac{1.5×3.2}{3}"


V=1.6L

So after adding 0.2 mole of gas new volume of gas is 1.6L